f(x)=4-2x
g(x)=square root of 3+4x
What is f o g?

Answer by Cromlix(4381)

Hi there,
f(x) = 4 - 2x
g(x) = √(3 + 4x)
fog = f(g(x))
f o g = 4 - 2(√(3 + 4x))
f o g = 4 - 2√3 - 8x
Hope this helps :-)